OpenStudy (anonymous):

Find the complete solution for the linear system if possible. x + y + z =8 x + 2y + 2z =14 x + y - z =2

OpenStudy (anonymous):

subtract the third equation from the first and post what you get.

OpenStudy (anonymous):

0+0-2z=-6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

right. so if -2z = -6, then z=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

ok good. substitute that in the first equation

OpenStudy (anonymous):

x+y+3=8?

OpenStudy (anonymous):

yes, so x+y =?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

okay now look at the 2nd equation x+2y+2z=14 you know that z=3 x+2y = 8 x+y+y=8 you know that x+y = 5 5+y =8

OpenStudy (anonymous):

so what would the final answer be written like?

OpenStudy (anonymous):

what is y?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

and therefore x =?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

x+y=5, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so plug in 3 for y giving you x =2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

right. you have your answer